# When Data is the Algorithm: A Systematic Study and Curation of Preference Optimization Datasets

### v1

@@ D.1 Dataset Compositions per Task Category
Tables 5 and 6 compare the composition of each DPO dataset by task category, as labeled by Magpie. This provides a sample-level view of how different task types are distributed across all examined open-source datasets, as well as our curated DPO mixtures.
@@ D.5 Language and Safety Analysis
Fig. 21 shows the distribution of language and safety across all examined open-source DPO datasets. We find that most datasets are predominantly English (96–99%), with the exception of TuluDPO, which includes approximately 12% non-English samples, primarily in French and German. All datasets are also rated as overwhelmingly safe (95–98%). As already discussed by tulutalk, language and safety are not significantly correlated with post-training performance. For this reason, we do not include these attributes in our data curation recipes. Nevertheless, we release all language and safety labels as part of our annotated datasets.
@@ F.1.1 Supervised Fine-Tuning (SFT)
Since Llama-3.1-8B [grattafiori2024llama3herdmodels] and Qwen-2.5-7B [yang2024qwen2] do not provide open SFT checkpoints, we fine-tune both models on the same SFT dataset to ensure consistency. Specifically, we use TuluSFT, a well-established, popular dataset curated by lambert2025tulu3pushingfrontiers for post-training Llama-3.1 models. Given that Llama and Qwen are of comparable size, we adopt the same hyperparameters for SFT tuning. The full training configurations are summarized in Table 8.
We note that by default, Open-Instruct applies a sum-reduction over token-level losses, rather than the more commonly used mean-reduction. This design choice ensures length-equitable weighting, where short and long sequences contribute proportionally to the total loss, preventing shorter examples from disproportionately influencing the gradient due to having fewer tokens. This leads to more stable optimization by avoiding fluctuations in loss scale caused by variation in batch composition or sequence length distributions. We refer to a more detailed discussion in lambert2025tulu3pushingfrontiers.
@@ F.1.2 Direct Preference Optimization (DPO)
Unlike SFT, where loss aggregation across tokens requires careful handling to avoid length bias, the DPO objective is inherently length-normalized [rafailov2024directpreferenceoptimizationlanguage, lambert2025tulu3pushingfrontiers]. This formulation ensures that completions of different lengths are treated equitably, and it avoids instabilities from batch composition or sequence length distributions observed in prior SFT pipelines.
To ensure consistency, we fix the training hyperparameters for each model across all datasets and apply DPO strictly on models that have already been fine-tuned via SFT using their instruction-tuning datasets. If available, we use the SFT checkpoints from HuggingFace. For Llama-3.1-8B-TuluSFT and Qwen-2.5-7B-TuluSFT, we adopt the same hyperparameters as in lambert2025tulu3pushingfrontiers. For the additional six open models, Apertus-8B-SFT [swissai2025apertus], OLMo-2-7B-SFT [olmo20242olmo2furious], SmolLM-3-3B-SFT [bakouch2025smollm3], Instella-3B-SFT [Instella], SmolLM-2-1.7B-SFT [allal2025smollm2smolgoesbig], and OLMo-2-1B-SFT [olmo20242olmo2furious], we adopt the hyperparameters provided in their respective papers. Furthermore, we use the length-normalized DPO loss (dpo_loss_type=norm), following the recommendation in lambert2025tulu3pushingfrontiers. Tables 9 and 10 report the DPO training configurations used for large and medium/small models, respectively.
@@ F.2 Evaluation Setup
We assess model performance using the LM Evaluation Harness framework [lm-eval-harness], a widely adopted standard for evaluating language models across diverse benchmark suites. To ensure a comprehensive and task-diverse evaluation, we include benchmarks from Open LLM Leaderboards V1 and V2 [leaderboardv1, open-llm-leaderboard-v2] to gauge downstream task performance under competitive public standards. This spans Knowledge (i.e., MMLU [hendryckstest2021], MMLU-Pro [wang2024mmluprorobustchallengingmultitask], TruthfulQA [lin-etal-2022-truthfulqa], GPQA [rein2023gpqagraduatelevelgoogleproofqa]), Reasoning (i.e., ARC-C [Clark2018ThinkYH], BBH [suzgun2022challenging], MuSR [sprague2024musrtestinglimitschainofthought]), Commonsense Understanding (i.e., HellaSwag [zellers2019hellaswag], WinoGrande [sakaguchi2019winogrande]), Instruction Following (i.e., IF-Eval [zhou2023instructionfollowing]), and Mathematical Reasoning (i.e., GSM8K [cobbe2021training], MATH [hendrycksmath2021]). In addition, we evaluate Coding via HumanEval and HumanEval+ [chen2021codex]. These benchmarks ensure a fair comparison between models and data mixtures.
@@ F.3 Additional DPO Results
To demonstrate the effectiveness of UltraMix across different architectures and scales, we evaluate six additional open models that have publicly released their corresponding SFT checkpoints: Apertus-8B-SFT [swissai2025apertus], OLMo-2-7B-SFT [olmo20242olmo2furious], SmolLM-3-3B-SFT [bakouch2025smollm3], Instella-3B-SFT [Instella]), SmolLM-2-1.7B-SFT [allal2025smollm2smolgoesbig], and OLMo-2-1B-SFT [olmo20242olmo2furious]. In contrast to our TuluSFT-tuned Llama and Qwen variants, this enables an important axis for evaluation with models that were SFT-tuned using different datasets.
Tables 11 to 16 report evaluation results across all considered benchmarks and DPO datasets, including our curated mixtures: UltraMix-170k (UM-170k), UltraMix-187k (UM-187k), and UltraMix-190k (UM-190k). Among the original DPO datasets, TuluDPO performs best, followed by UltraFeedback and ORPO. Our curated UltraMix-190k mixture consistently outperforms TuluDPO, achieving substantial improvements on nearly all benchmarks while being 30% smaller in size. These additional results corroborate the findings presented in the main paper and further demonstrate the effectiveness of our curated UltraMix dataset across diverse model architectures and scales.
@@ F.3.6 OLMo-2-1B-SFT
F.4 DPO Results without Preference Reward-Based Filtering
To demonstrate that coherent preference rewards are essential for constructing effective DPO data mixtures, we report evaluation results for UltraMix-No-Preference-Reward-Filter (UM-No-PF) in Table 17, a variant of UltraMix where we omitted the preference reward-based filtering step and only applied quality- and task-aware filtering according to our principled curation recipe in Fig. 23. The results show that despite careful quality-based filtering and task-aware addition of instruction following data, UM-No-RF underperforms both TuluDPO and our reward-filtered UltraMix variants UM-170k and UM-190k in overall and OpenLLM leaderboard scores. This suggest that DPO training is particularly sensitive to the preservation of coherent preference orderings, which implies that relying solely on quality and task-based filtering may introduce distortions that obscure preference consistency and degrade downstream performance.
This important insight distinguishes our work from other post-training data curation recipes, such as for SFT [tulutalk], showing that DPO datasets require more than careful quality- and task-based curation, and rely on the preservation of a clear and consistent preference ordering. The preference reward filtering step in our principled curation recipe removes incoherent preference pairs (where a chosen completion receives a lower reward than its rejected counterpart), thereby improving consistency in preference assignment and allowing for subsequent quality- and task-based filtering.
@@ F.4 DPO Results without Preference Reward-Based Filtering
F.5 Efficiency Gains

### v2

@@ D.1 Dataset Compositions per Task Category
Tables 7 and 8 compare the composition of each DPO dataset by task category, as labeled by Magpie. This provides a sample-level view of how different task types are distributed across all examined open-source datasets, as well as our curated DPO mixtures.
@@ D.5 Language and Safety Analysis
Fig. 21 shows the distribution of language and safety across all examined open-source DPO datasets. We find that most datasets are predominantly English (96–99%), with the exception of TuluDPO, which includes approximately 12% non-English samples, primarily in French and German. All datasets are also rated as overwhelmingly safe (95–98%). As already discussed by Djuhera et al. [2025], language and safety are not significantly correlated with post-training performance. For this reason, we do not include these attributes in our data curation recipes. Nevertheless, we release all language and safety labels as part of our annotated datasets.
@@ F.1.1 Supervised Fine-Tuning (SFT)
Since Llama-3.1-8B [Grattafiori et al., 2024] and Qwen-2.5-7B [Yang et al., 2025] do not provide open SFT checkpoints, we fine-tune both models on the same SFT dataset to ensure consistency. Specifically, we use TuluSFT, a well-established, popular dataset curated by Lambert et al. [2025] for post-training Llama-3.1 models. Given that Llama and Qwen are of comparable size, we adopt the same hyperparameters for SFT tuning. The full training configurations are summarized in Table 10.
We note that by default, Open-Instruct applies a sum-reduction over token-level losses, rather than the more commonly used mean-reduction. This design choice ensures length-equitable weighting, where short and long sequences contribute proportionally to the total loss, preventing shorter examples from disproportionately influencing the gradient due to having fewer tokens. This leads to more stable optimization by avoiding fluctuations in loss scale caused by variation in batch composition or sequence length distributions. We refer to a more detailed discussion in Lambert et al. [2025].
@@ F.1.2 Direct Preference Optimization (DPO)
Unlike SFT, where loss aggregation across tokens requires careful handling to avoid length bias, the DPO objective is inherently length-normalized [Rafailov et al., 2023, Lambert et al., 2025]. This formulation ensures that completions of different lengths are treated equitably, and it avoids instabilities from batch composition or sequence length distributions observed in prior SFT pipelines.
To ensure consistency, we fix the training hyperparameters for each model across all datasets and apply DPO strictly on models that have already been fine-tuned via SFT using their instruction-tuning datasets. If available, we use the SFT checkpoints from HuggingFace. For Llama-3.1-8B-TuluSFT and Qwen-2.5-7B-TuluSFT, we adopt the same hyperparameters as in Lambert et al. [2025]. For the additional six open models, Apertus-8B-SFT [Apertus-Team, 2025], OLMo-2-7B-SFT [OLMo et al., 2024], SmolLM-3-3B-SFT [Bakouch et al., 2025], Instella-3B-SFT [Liu et al., 2025], SmolLM-2-1.7B-SFT [Allal et al., 2025], and OLMo-2-1B-SFT [OLMo et al., 2024], we adopt the hyperparameters provided in their respective papers. Furthermore, we use the length-normalized DPO loss (dpo_loss_type=norm), following the recommendation in Lambert et al. [2025]. Tables 11 and 12 report the DPO training configurations used for large and medium/small models, respectively.
@@ F.2 Evaluation Setup
We assess model performance using the LM Evaluation Harness framework [Gao et al., 2024], a widely adopted standard for evaluating language models across diverse benchmark suites. To ensure a comprehensive and task-diverse evaluation, we include benchmarks from Open LLM Leaderboards V1 and V2 [Fourrier et al., 2023, 2024] to gauge downstream task performance under competitive public standards. This spans Knowledge (i.e., MMLU [Hendrycks et al., 2021a], MMLU-Pro [Wang et al., 2024b], TruthfulQA [Lin et al., 2022], GPQA [Rein et al., 2023]), Reasoning (i.e., ARC-C [Clark et al., 2018], BBH [Suzgun et al., 2022], MuSR [Sprague et al., 2024]), Commonsense Understanding (i.e., HellaSwag [Zellers et al., 2019], WinoGrande [Sakaguchi et al., 2019]), Instruction Following (i.e., IF-Eval [Zhou et al., 2023]), and Mathematical Reasoning (i.e., GSM8K [Cobbe et al., 2021], MATH [Hendrycks et al., 2021b]). In addition, we evaluate Coding via HumanEval and HumanEval+ [Chen et al., 2021]. These benchmarks ensure a fair comparison between models and data mixtures.
@@ F.3 Additional DPO Results
To demonstrate the effectiveness of UltraMix across different architectures and scales, we evaluate six additional open models that have publicly released their corresponding SFT checkpoints: Apertus-8B-SFT [Apertus-Team, 2025], OLMo-2-7B-SFT [OLMo et al., 2024], SmolLM-3-3B-SFT [Bakouch et al., 2025], Instella-3B-SFT [Liu et al., 2025]), SmolLM-2-1.7B-SFT [Allal et al., 2025], and OLMo-2-1B-SFT [OLMo et al., 2024]. In contrast to our TuluSFT-tuned Llama and Qwen variants, this enables an important axis for evaluation with models that were SFT-tuned using different datasets.
Tables 13 to 18 report evaluation results across all considered benchmarks and DPO datasets, including our curated mixtures: UltraMix-170k (UM-170k), UltraMix-187k (UM-187k), and UltraMix-190k (UM-190k). Among the original DPO datasets, TuluDPO performs best, followed by UltraFeedback and ORPO. Our curated UltraMix-190k mixture consistently outperforms TuluDPO, achieving substantial improvements on nearly all benchmarks while being 30% smaller in size. These additional results corroborate the findings presented in the main paper and further demonstrate the effectiveness of our curated UltraMix dataset across diverse model architectures and scales.
@@ F.3.6 OLMo-2-1B-SFT
F.4 Assessing the Role of Individual Filtering Steps
To understand the contribution of each filtering step in our data-curation recipe, we provide granular ablations and show that UltraMix’s effectiveness arises from the combination of diverse signals, rather than from coherence reward or quality-based filtering alone.
We first demonstrate that coherent preference rewards are essential for constructing effective DPO data mixtures and report evaluation results for UltraMix-No-Preference-Reward-Filter (UM-No-PF) in Table 19, a variant of UltraMix where we omitted the preference reward-based filtering step and only applied quality- and task-aware filtering according to our principled curation recipe in Fig. 23. The results show that despite careful quality-based filtering and task-aware addition of instruction following data, UM-No-PF underperforms both TuluDPO and our reward-filtered UltraMix variants UM-170k and UM-190k in overall and OpenLLM leaderboard scores. This suggest that DPO training is particularly sensitive to the preservation of coherent preference orderings, which implies that relying solely on quality and task-based filtering may introduce distortions that obscure preference consistency and degrade downstream performance.
To further substantiate these claims, we provide additional evaluations in Table 20 for (a) UM-only-PF, which applies preference-reward filtering only, and (b) UM-only-QF, which applies input-quality filtering only, and compare them against (c) UM-No-PF and (d) the full UltraMix recipe that combines all steps. The results show that no single step outperforms the TuluDPO baseline on its own. UM-only-QF and UM-No-PF generally outperform UM-only-PF, showing preference filtering is important but not the main driver. Thus, best performance is only achieved when every step is combined. These ablations clarify the role of each step in our curation recipe, demonstrating that all stages are complementary and the effectiveness of UltraMix stems from the combination of signals.
This important insight distinguishes our work from other post-training data curation recipes, such as for SFT [Djuhera et al., 2025], showing that DPO datasets require a joint curation that involves careful quality- and task-based curation, and relies on the preservation of a clear and consistent preference ordering. The preference reward filtering step in our principled curation recipe removes incoherent preference pairs (where a chosen completion receives a lower reward than its rejected counterpart), thereby improving consistency in preference assignment and allowing for subsequent quality- and task-based filtering.
@@ F.4 Assessing the Role of Individual Filtering Steps
Table 20: DPO results for Llama-3.1-8B-TuluSFT and Qwen-2.5-7B-TuluSFT trained on ablated DPO mixtures UM-only-PF, UM-only-QF, and UM-No-PF, compared to TuluDPO on Open LLM Leaderboards (averaged) and code benchmarks. The overall average is across all benchmarks. Best scores are in bold.
Llama-3.1-8B-TuluSFT Qwen-2.5-7B-TuluSFT
Benchmark TuluDPO UM-only-PF UM-only-QF UM-No-PF UM-190k TuluDPO UM-only-PF UM-only-QF UM-No-PF UM-190k
Knowledge
MMLU (5-shot) 63.47 63.42 63.35 62.69 64.61 73.10 72.90 73.17 73.05 74.01
MMLU-Pro (5-shot) 28.98 29.58 28.79 28.55 30.96 43.48 42.56 43.43 43.52 44.65
TruthfulQA (0-shot) 56.78 57.43 58.44 56.80 63.32 54.46 55.68 54.13 55.20 58.00
GPQA (0-shot) 29.61 30.21 29.78 29.52 31.87 31.29 30.62 29.95 31.12 33.03
Reasoning
ARC-C (25-shot) 57.93 57.68 58.02 57.76 58.70 60.32 61.97 61.18 60.92 62.43
BBH (3-shot) 40.46 41.22 40.90 40.28 44.96 48.08 49.78 48.43 48.29 51.76
MuSR (0-shot) 41.93 40.08 42.06 41.87 44.02 47.16 44.12 46.43 47.30 48.82
Commonsense
HellaSwag (10-shot) 64.85 62.26 63.94 62.84 64.82 62.70 62.07 63.03 62.16 63.89
WinoGrande (5-shot) 75.30 74.03 76.56 75.72 77.06 72.96 70.61 71.11 72.19 74.64
Instruction Following
IF-Eval (0-shot) 80.35 77.61 79.74 79.93 81.13 78.04 76.55 77.70 75.39 79.88
Math
GSM8K (5-shot) 79.48 74.97 76.65 79.76 82.48 76.84 70.13 71.49 77.69 82.70
MATH (4-shot) 22.66 13.82 20.17 19.79 23.56 43.13 40.30 42.60 42.88 49.55
Code (pass@1)
HumanEval 67.24 59.05 64.07 60.98 69.05 80.49 73.66 73.78 74.39 82.27
HumanEval+ 46.36 43.07 43.61 43.59 48.08 61.83 57.89 58.05 59.03 63.05
Leaderboards
Open LLM Leaderboard 1 66.30 64.96 66.16 65.93 68.50 66.73 65.56 65.68 66.87 69.28
Open LLM Leaderboard 2 40.66 38.75 40.24 39.99 42.75 48.53 47.32 48.09 48.08 51.28
Overall 53.96 51.74 53.29 52.86 56.04 59.56 57.77 58.18 58.79 62.05
F.5 Assessing the Role of Incoherent Preference Orders
As discussed in the main body, our independently trained reward model disagrees with the preference orderings in many of the existing DPO datasets we evaluate. There are several likely reasons for this mismatch. For instance, the use of GPT-4-generated scalar scores in UltraFeedback [Cui et al., 2024] does not guarantee consistent or reliable preference signals, and the authors explain why:
  • Ambiguous Scores: GPT-4 frequently assigns identical scores to chosen/rejected responses as completions are near-ties (sourced from 17 LLMs).
  • Limited Human Agreement: UltraFeedback reports only 59–68% human agreement, which is close to the  75% consistency we observe.
  • Reward Model Accuracy: their own reward model achieves only 66–71% accuracy, confirming nontrivial noise in its preference signals.
We provide additional empirical evidence for this point: Tables 21 and 22 show that, for both Llama and Qwen models, filtering out misaligned preference pairs (“[Pref]" variants) consistently improves downstream performance on almost all benchmarks, especially on aggregate leaderboard metrics.
Table 21: DPO results for Llama-3.1-8B-TuluSFT trained on all preference datasets with (indicated by [Pref]) and without reward-based filtering. The overall average is across all benchmarks. Best scores are in bold.
Llama-3.1-8B-TuluSFT
Benchmark SFT TuluDPO TuluDPO [Pref] UltraFB UltraFB [Pref] ORPO ORPO [Pref] HelpSteer HelpSteer [Pref] CodePref CodePref [Pref]
Knowledge
MMLU (5-shot) 62.30 63.47 64.14 62.53 62.96 62.31 62.96 62.04 62.77 59.96 60.02
MMLU-Pro (5-shot) 28.08 28.98 28.96 28.41 28.06 27.90 27.98 27.43 27.31 27.53 27.69
TruthfulQA (0-shot) 46.84 56.78 61.31 50.26 54.77 52.26 56.88 47.43 48.22 56.15 57.35
GPQA (0-shot) 28.44 29.61 29.69 28.69 29.61 29.70 29.78 29.36 30.70 29.95 30.03
Reasoning
ARC-C (25-shot) 54.95 57.93 57.40 56.91 57.08 56.91 56.23 54.52 54.61 45.05 45.54
BBH (3-shot) 38.59 40.46 43.93 39.91 41.61 41.80 42.13 38.15 39.48 36.07 36.49
MuSR (0-shot) 43.25 41.93 42.68 41.93 41.67 43.78 43.27 42.06 42.14 41.53 42.74
Commonsense
HellaSwag (10-shot) 60.41 64.85 64.24 62.29 62.83 60.08 60.78 61.20 61.39 53.38 53.46
WinoGrande (5-shot) 76.40 75.30 76.19 75.53 76.40 74.27 75.69 76.16 76.32 69.93 70.75
Instruction Following
IF-Eval (0-shot) 72.45 80.35 80.74 71.59 72.18 71.65 73.12 73.93 74.94 69.90 72.23
Math
GSM8K (5-shot) 76.19 79.48 80.74 78.47 80.06 81.96 81.83 76.88 78.32 76.35 76.44
MATH (4-shot) 12.08 22.66 22.79 18.81 20.02 20.39 21.98 14.88 16.16 15.11 15.33
Code (pass@1)
HumanEval 57.93 67.24 68.76 61.59 68.29 64.63 65.63 59.15 61.91 70.68 70.24
HumanEval+ 43.29 46.36 47.56 42.49 44.39 41.63 43.11 42.93 44.76 50.61 50.41
Leaderboards
Open LLM Leaderboard 1 62.85 66.30 67.34 64.33 65.68 64.63 65.73 63.04 63.60 60.14 60.59
Open LLM Leaderboard 2 37.15 40.66 41.46 38.22 38.86 39.20 39.71 37.64 38.46 36.68 37.42
Overall 50.09 53.96 54.94 51.39 52.85 52.09 52.96 50.44 51.36 50.16 50.62
Table 22: DPO results for Qwen-2.5-7B-TuluSFT trained on all preference datasets with (indicated by [Pref]) and without reward-based filtering. The overall average is across all benchmarks. Best scores are in bold.
Qwen-2.5-7B-TuluSFT
Benchmark SFT TuluDPO TuluDPO [Pref] UltraFB UltraFB [Pref] ORPO ORPO [Pref] HelpSteer HelpSteer [Pref] CodePref CodePref [Pref]
Knowledge
MMLU (5-shot) 72.41 73.10 73.28 73.32 73.47 73.12 73.36 72.38 73.38 72.82 72.85
MMLU-Pro (5-shot) 43.32 43.48 43.40 43.94 43.88 43.73 43.80 43.07 44.11 43.38 43.54
TruthfulQA (0-shot) 51.64 54.46 55.30 53.95 54.49 53.53 53.62 52.27 52.51 52.10 52.39
GPQA (0-shot) 30.96 31.29 31.51 31.04 31.12 30.61 30.89 30.05 31.12 30.37 31.35
Reasoning
ARC-C (25-shot) 59.22 60.32 61.35 60.32 60.41 60.07 60.41 59.64 59.73 59.56 59.96
BBH (3-shot) 47.93 48.08 50.03 48.57 49.71 47.14 48.00 47.19 48.38 47.72 48.42
MuSR (0-shot) 48.15 47.16 47.71 47.75 47.92 47.88 48.09 48.41 48.85 46.30 47.09
Commonsense
HellaSwag (10-shot) 60.85 62.70 62.64 61.80 61.84 61.07 61.40 61.12 61.22 61.22 62.27
WinoGrande (5-shot) 73.17 72.96 73.19 72.85 73.07 72.06 72.51 72.11 74.27 70.48 70.40
Instruction Following
IF-Eval (0-shot) 68.06 78.04 78.30 75.74 76.06 73.02 74.35 70.16 72.80 71.05 73.19
Math
GSM8K (5-shot) 71.27 76.84 77.70 74.93 75.54 77.33 78.18 71.87 72.86 72.24 73.61
MATH (4-shot) 36.21 43.13 49.38 39.19 41.06 41.92 42.84 37.72 39.79 36.34 39.04
Code (pass@1)
HumanEval 72.66 80.49 80.88 76.05 78.66 78.66 78.89 76.27 79.27 84.54 84.13
HumanEval+ 55.85 61.83 61.98 58.01 59.27 59.10 60.98 56.02 61.59 65.49 65.27
Leaderboards
Open LLM Leaderboard 1 64.76 66.73 67.24 66.19 66.47 66.20 66.58 64.90 65.66 64.74 65.25
Open LLM Leaderboard 2 45.77 48.53 50.06 47.70 48.29 47.38 48.00 46.10 47.51 45.86 47.10
Overall 56.55 59.56 60.48 58.39 59.04 58.52 59.09 57.02 58.56 58.11 58.82
F.6 Efficiency Gains